# Manifold-aware Representation Learning for Degradation-agnostic Image Restoration

### v1

@@ 5 Experiments
Zero-Shot Cross-Domain (i.e., Underwater) Setting. We evaluate our method’s generalization under a challenging zero-shot setting using real-world underwater images. As shown in Tab. 6, MIRAGE -S achieves 17.29 dB and 0.773 SSIM, surpassing the best prior method MoCE-IR zamfir2025moce by a clear margin (+1.38 dB PSNR), while being more compact. Importantly, our model has never seen underwater data during training. This demonstrates that our adaptive modeling not only fits mixed degradations but also transfers robustly to unseen, conditions.
Efficiency Comparison. Tab. 6 presents a detailed comparison of PSNR, memory usage, parameter count, and FLOPs. Our Tiny model (MIRAGE -T) achieves the best efficiency-performance trade-off: with only 6.21M parameters and 16G FLOPs, it outperforms all prior methods, including larger models like PromptIR potlapalli2023promptir and MoCE-IR-S zamfir2025moce . Notably, MIRAGE -T surpasses MoCE-IR-S by +0.26 dB while requiring less than half the computational cost. Even our Small variant (MIRAGE -S) exceeds full MoCE-IR in both PSNR (+0.18 dB) and FLOPs (27G vs. 75G). These results validate that our design achieves strong restoration quality without sacrificing computational efficiency.
Visual Comparison. Across various restoration tasks, MIRAGE consistently produces sharper structures, richer textures, and fewer artifacts compared to existing methods. Fig. 1 demonstrates its strong generalization and ability to recover fine-grained details under diverse degradations, showing superior performance than others. More visual results are provided in our Supp. Mat.
5.2 Ablation Study.
We conduct ablation studies to assess the contribution of key components in MIRAGE , as summarized in Tab. 7. Starting from an attention-only baseline (32.23 dB, 19.89M), we progressively integrate each module while reducing overall complexity. Removing the dynamic convolution branch (w/o DynamicConv) causes a 0.56 dB drop, indicating its importance for local spatial modeling.
Table 7: Ablation Study of MIRAGE -T on 3 Degradation Setting.
Ablaton Parms. Results
@@
The channel-wise MLP (w/o C-MLP) also plays a critical role, with a 0.38 dB performance loss. Replacing gated fusion with naive concatenation (w/o Fusion) leads to a further 0.20 dB drop, confirming that explicit feature integration is more effective. On the regularization side, removing contrastive learning (w/o CL & SPD) or only the SPD module degrades performance by 0.14 dB and 0.24 dB respectively, highlighting the benefit of manifold-aware cross-depth alignment. Overall, each component contributes to the final performance. Our full model achieves the best balance between accuracy and efficiency with only 6.21M parameters and 32.77 dB PSNR.
@@ 6 Conclusion
In this paper, we present MIRAGE , a unified and lightweight framework for all-in-one image restoration. By decomposing features into attention-, convolution-, and MLP-based branches, MIRAGE captures global context, local textures, and channel-wise statistics in a complementary and efficient manner. To bridge the semantic gap across depths, we introduce a shallow-latent contrastive learning scheme that aligns early and latent representations via second-order feature statistics on the SPD manifold. This improves cross-stage consistency and enhances representation discriminability without additional inference cost. Extensive experiments across diverse degradations and cross-domain settings show that MIRAGE achieves state-of-the-art performance with minimal parameters and FLOPs. We believe its design offers a scalable path for lightweight, adaptive restoration, with potential extensions to video and multi-modal scenarios. Please refer to our Supp. Mat. for additional discussions, implementation details, and visual results.
@@ A.1 Datasets
3 Degradation Datasets. For both the All-in-One and single-task settings, we follow the evaluation protocols established in prior works li2022all ; potlapalli2023promptir ; zamfir2025moce , utilizing the following datasets: For image denoising in the single-task setting, we combine the BSD400 arbelaez2010contour and WED ma2016waterloo datasets, and corrupt the images with Gaussian noise at levels $\sigma\in\{15,25,50\}$. BSD400 contains 400 training images, while WED includes 4,744 images. We evaluate the denoising performance on BSD68 martin2001database and Urban100 huang2015single . For single-task deraining, we use Rain100L yang2020learning , which provides 200 clean/rainy image pairs for training and 100 pairs for testing. For single-task dehazing, we adopt the SOTS dataset li2018benchmarking , consisting of 72,135 training images and 500 testing images. Under the All-in-One setting, we train a unified model on the combined set of the aforementioned training datasets for 120 epochs and directly test it across all three restoration tasks.
@@
5 Degradation Datasets. The 5-degradation setting is built upon the 3-degradation setting, with two additional tasks included: deblurring and low-light enhancement. For deblurring, we adopt the GoPro dataset nah2017deep , which contains 2,103 training images and 1,111 testing images. For low-light enhancement, we use the LOL-v1 dataset wei2018deep , consisting of 485 training images and 15 testing images. Note that for the denoising task under the 5-degradation setting, we report results using Gaussian noise with $\sigma=25$. The training takes 130 epochs.
@@
Composited Degradation Datasets. Regarding the composite degradation setting, we use the CDD11 dataset guo2024onerestore . CDD11 consists of 1,183 training images for: (i) 4 kinds of single-degradation types: haze (H), low-light (L), rain (R), and snow (S); (ii) 5 kinds of double-degradation types: low-light + haze (l+h), low-light+rain (L+R), low-light + snow (L+S), haze + rain (H+R), and haze + snow (H+S). (iii) 2 kinds of Triple-degradation type: low-light + haze + rain (L+H+R), and low-light + haze + snow (L+H+S). We train our method for 150 epochs (fewer than 200 epochs than MoCE-IR zamfir2025moce ), and we keep all other settings unchanged.
@@
Adverse Weather Removal Datasets. For the deweathering tasks, we follow the experimental setups used in TransWeather Transweather and WGWSNet ZhuWFYGDQH23 , evaluating the performance of our approach on multiple synthetic datasets. We assess the capability of MIRAGE across three challenging tasks: snow removal, rain streak and fog removal, and raindrop removal. The training set, referred to as “AllWeather”, is composed of images from the Snow100K liu2018desnownet , Raindrop qian2018attentive , and Outdoor-Rain li2019heavy datasets. For testing, we evaluate our model on the following subsets: Snow100K-S (16,611 images), Snow100K-L (16,801 images), Outdoor-Rain (750 images), and Raindrop (249 images). Same as Histoformer Histoformer_SunRGWC24 , we train MIRAGE on “AllWeather” with 300,000 iterations.
@@
Zero-Shot Underwater Image Enhancement Dataset. For the zero-shot underwater image enhancement setting, we follow the evaluation protocol of DCPT hu2025universal by directly applying our model, trained under the 5-degradation setting, on the UIEB dataset li2019underwater without any finetuning. UIEB consists of two subsets: 890 raw underwater images with corresponding high-quality reference images, and 60 challenging underwater images. We evaluate our zero-shot performance on the 890-image subset with available reference images.
@@ A.2 Implementation Details
[…]
@@ A.3 Optimization Objectives
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{1}}+\lambda_{fre}\times\mathcal{L}_{\text{Fourier}}+\lambda_{ctrs}\times\mathcal{L}_{\text{SPD}}.$ (6)
@@
Specifically, we adopt an $\ell_{1}$ loss that adopted in IR tasks potlapalli2023promptir ; zamfir2025moce ; li2022all ; cui2025adair ; ren2024sharing , defined as $\mathcal{L}_{\text{1}}=|\hat{x}-x|_{1}$, to enforce pixel-wise similarity between the restored image $\hat{x}$ and the ground-truth image $x$. $\mathcal{L}_{\text{Fourier}}$, as utilized in MoCE-IR zamfir2025moce ; cui2025adair , to enhance frequency-domain consistency, the real-valued Fourier loss, is defined as:
$\mathcal{L}_{\text{Fourier}}=\left\|\mathcal{F}_{\text{real}}(\hat{x})-\mathcal{F}_{\text{real}}(x)\right\|_{1}+\left\|\mathcal{F}_{\text{imag}}(\hat{x})-\mathcal{F}_{\text{imag}}(x)\right\|_{1},$ (7)
[…]
Appendix B More Method Details & Supplementary Experiments
B.1 1 Deg. Comparison
@@ Appendix B More Method Details & Supplementary Experiments
DehazeNet cai2016dehazenet - 22.46 .851
MSCNN ren2016single - 22.06 .908
AODNet li2017aod - 20.29 .877
EPDN  qu2019enhanced - 22.57 .863
FDGAN dong2020fdgan - 23.15 .921
AirNet li2022all 9M 23.18 .900
PromptIR potlapalli2023promptir 36M 31.31 .973
@@
DIDMDN zhang2018density - 23.79 .773
UMR yasarla2019uncertainty - 32.39 .921
SIRR wei2019semi - 32.37 .926
MSPFN jiang2020multi - 33.50 .948
LPNet gao2019dynamic - 23.15 .921
AirNet li2022all 9M 34.90 .977
PromptIR potlapalli2023promptir 36M 37.04 .979
@@
DnCNN zhang2017beyond - 33.89 .930 31.23 .883 27.92 .789
IRCNN  zhang2017learning - 33.87 .929 31.18 .882 27.88 .790
FFDNet  zhang2018ffdnet - 33.87 .929 31.21 .882 27.96 .789
BRDNet tian2000brdnet - 34.10 .929 31.43 .885 28.16 .794
AirNet li2022all 9M 34.14 .936 31.48 .893 28.23 .806
PromptIR potlapalli2023promptir 36M 34.34 .938 31.71 .897 28.49 .813
PromptIR potlapalli2023promptir (Repdoduce) 36M 34.15 .934 31.50 .894 28.33 .807
@@
Single-Degradation. Tab. B presents the results of MIRAGE trained individually for dehazing, deraining, and denoising. Across all tasks, our method consistently surpasses previous state-of-the-art approaches, including PromptIR potlapalli2023promptir and its reproduced variant. Particularly in the denoising task on BSD68, our model achieves the best performance across all noise levels. Interestingly, while the single-task models perform competitively, they are slightly outperformed by the all-in-one version in dehazing and deraining. This observation suggests that different degradation types may exhibit shared structures, and that learning them jointly can lead to more generalizable representations.
@@
1:$\alpha\in\mathbb{R}^{B\times C\times H\times W}$triangleright Input feature map
2:$\alpha^{\prime}\in\mathbb{R}^{B\times C\times H\times W}$triangleright Output after dynamic depthwise conv
@@
4:$K\leftarrow\texttt{AdaptiveAvgPool2D}(\alpha)$triangleright Global context pooling
5:$K\leftarrow\texttt{Conv2D}(K,\ 1\times 1,\ \text{out\_ch}=C)$triangleright Linear projection
6:$K\leftarrow\texttt{GELU}(K)$triangleright Non-linear activation
7:$K\leftarrow\texttt{Conv2D}(K,\ 1\times 1,\ \text{out\_ch}=C\cdot k^{2})$triangleright Generate kernel weights
8:$K\leftarrow\texttt{Reshape}(K,\ [B\cdot C,\ 1,\ k,\ k])$triangleright Form depthwise filters
@@
10:$\alpha_{\text{flat}}\leftarrow\texttt{Reshape}(\alpha,\ [1,\ B\cdot C,\ H,\ W])$triangleright Prepare for grouped conv
@@
12:$\alpha^{\prime}\leftarrow\texttt{Reshape}(\alpha^{\prime}_{\text{flat}},\ [B,\ C,\ H,\ W])$triangleright Reshape back to original shape
@@
B.2 Details of the Design for the proposed Mixed Backbone.
@@
# $patch\_embedding$: shallow convolutional patch embedding
# $refinement\_conv$: the refinement block and the final convolution
# $spd$: compute SPD feature
for ${x}$ in loader: # load a minibatch x with n samples
$F_{\text{shallow}}$ = $patch\_embedding$(${x}$) # Convolutional Patch Embedding
@@
$\mathbf{C}_{s}$, $\mathbf{C}_{l}$= $spd$($F_{\text{shallow}}$), $spd$($F_{\text{latent}}$) # Compurte SPD (Symmetric Positive Definite) manifold features
$\mathbf{z}_{s}$, $\mathbf{z}_{l}$ = $proj\_norm$($\mathbf{C}_{s}$), $proj\_norm$($\mathbf{C}_{l}$) # Projection and normalize
@@
$\hat{x}$ = $refinement\_conv$($F_{\text{recon}}$)
L = $\mathcal{L}_{\text{1}}$(${x}$, $\hat{x}$) + $\lambda_{fre}\times$$\mathcal{L}_{\text{Fourier}}$ (${x}$, $\hat{x}$) + $\lambda_{ctrs}\times$$\mathcal{L}_{\text{SPD}}$($\mathbf{z}_{s}$, $\mathbf{z}_{l}$) # total loss
@@
update($f_{en}$, $f_{de}$, $patch\_embedding$, $refinement\_conv$) # SGD update
@@
def $\mathcal{L}_{\text{SPD}}$(a, b): # Real-valued Fourier loss
␣␣␣␣RefertoEq.5ofourmainmanuscript.
@@
B.3 Details of the Proposed SPD Contrastive Learning.
@@
B.4 Ablation Regarding the Optimization Objectives

### v2

@@ 5 Experiments
Zero-Shot Setting. We evaluate our method’s generalization under a challenging zero-shot setting with real-world underwater images. As shown in Tab. 6, MIRAGE -S achieves 17.29 dB and 0.773 SSIM, surpassing MoCE-IR (Zamfir et al., 2025) by +1.38dB PSNR, while being more compact. Importantly, our model never sees underwater data during training, yet our adaptive modeling not only fits mixed degradations but also transfers robustly to unseen conditions. Besides, we also followed the same experimental setting introduced by UniRestore Chen et al. (2025a) for the generalization ability evaluation. Meanwhile, the real-world evaluation presented in Tab. D shows that MIRAGE generalizes reliably to real-world, camera-captured degradations.
Efficiency Comparison. Tab. 6 compares PSNR, memory, parameters, and FLOPs. Our Tiny model (MIRAGE -T), with only 6.21M parameters and 16G FLOPs, delivers the best efficiency–performance trade-off, outperforming all prior methods, including larger models like PromptIR (Potlapalli et al., 2024) and MoCE-IR-S (Zamfir et al., 2025). It surpasses MoCE-IR-S by +0.26 dB while using less than half the computation, and even our Small variant (MIRAGE -S) exceeds full MoCE-IR in both PSNR (+0.18dB) and FLOPs (27G vs. 75G). These results confirm that our design achieves strong restoration quality without compromising efficiency.
Visual Comparison. MIRAGE effectively restores fine structural details and reliably suppresses subtle visual artifacts across diverse and unseen degradations (Fig. 1 and appendix).
5.2 Ablation Analysis & Discussion
Components ablation. Tab. 7 shows starting from an attention-only setting (32.23 dB, 19.89M), we progressively integrate each module while reducing complexity.
Table 7: Ablation Study of MIRAGE -T under the 3-Degradation Setting with Tiny model.
Ablaton Params. Results
@@ 5.2 Ablation Analysis & Discussion
Removing the dynamic convolution branch (w/o DynamicConv) causes a 0.56 dB drop, indicating its importance for local spatial modeling. The channel-wise MLP (w/o C-MLP) also plays a critical role, with a 0.38 dB performance loss. Naive concatenation (w/o Fusion) leads to a further 0.20 dB drop, confirming that explicit feature integration is more effective. On the regularization side, removing contrastive learning (w/o CL & SPD) or replacing SPD with Euclidean alignment degrades performance by 0.14 dB and 0.24 dB, indicating that structure-agnostic contrastive learning can misguide optimization, while manifold-aware alignment provides consistent benefits. Overall, each component contributes to the final performance. Our full model offers the best balance between accuracy and efficiency with only 6.21M parameters and 32.77 dB PSNR.
Figure 5: Shallow–latent cosine similarity across degradations. Contrastive alignment improves feature correlation.
Why shallow–latent Contrastive Alignment Matters. Different degradations rely on different feature levels: denoising and deraining benefit from shallow, texture-rich features, while dehazing and low-light enhancement require deeper semantic features; deblurring needs both. This heterogeneity makes unified modeling challenging. We therefore introduce contrastive alignment between shallow and latent stages to encourage semantic coordination. When shallow features dominate (e.g., denoising), alignment guides latent features to be more task-relevant; when latent features dominate (e.g., dehazing), shallow features inherit semantic consistency (Bertasius et al., 2015). Fig. 5 validates that contrastive alignment improves shallow–latent correlation, validating its necessity for cross-degradation generalization.
Figure 6: Shallow–latent similarity under three settings: (a) w/o CL (unstructured; off-diag $0.0784$, ratio $0.148$), (b) Euclidean CL (collapsed; off-diag $\approx 0.0024$, ratio $0.99$), (c) SPD CL (preserved; off-diag $0.0787$, ratio $0.149$).
Why Euclidean Fails and Why SPD Works? (Deraining Case Study) Euclidean contrastive learning collapses shallow–latent alignment by enforcing indiscriminate similarity, reducing both diagonal and off-diagonal terms to trivial constants, and erasing task cues. SPD, by aligning covariance matrices on a Riemannian manifold, preserves second-order dependencies and guides updates along meaningful directions. In the deraining case (Figure 6), Euclidean CL degenerates into near-constant similarity (off-diag $0.00237$, ratio $0.99$), while SPD maintains diagonal dominance and non-trivial off-diagonal structure ($0.0787$, ratio $0.149$), producing coherent patterns.
@@ 6 Conclusion
We presented MIRAGE , an efficient framework for degradation-agnostic image restoration that achieves a favorable balance between robustness and efficiency. Through channel-wise functional decomposition, the model repurposes redundant capacity into convolution-, attention-, and MLP-based branches, enabling complementary modeling of local textures, global context, and channel-wise statistics. To further enhance cross-degradation generalization, we introduced manifold regularization, aligning shallow and latent features in the SPD manifold space for more consistent and discriminative representations. Extensive experiments across diverse degradations, including mixed and unseen scenarios, demonstrate that MIRAGE achieves state-of-the-art performance. Inspired by the metaphor of a mirage, i.e., revealing the hidden reality beneath visual distortions, our framework learns degradation-agnostic representations by balancing global, local, and channel-wise information, providing a scalable foundation for future research in degradation-agnostic IR.
Acknowledgments
This work was partially supported by the FIS project GUIDANCE (Debugging Computer Vision Models via Controlled Cross-modal Generation) (No. FIS2023-03251).
Ethics Statement
Our work focuses on general-purpose image restoration, aiming to improve efficiency and robustness across diverse degradation types. The intended positive impact includes deployment in low-resource or safety-critical scenarios such as mobile photography, remote sensing, medical imaging, and environmental monitoring. At the same time, we recognize that improved restoration techniques could be misused for deceptive content editing or large-scale surveillance. We encourage responsible use of our method and provide our models and code with appropriate licenses and documentation to support transparency and ethical adoption. No personally identifiable or sensitive data were used in this research.
Reproducibility Statement
We aim to ensure reproducibility and transparency of our results. The MIRAGE framework is implemented in PyTorch with standard training protocols and evaluation metrics. Detailed descriptions of the architecture, training settings, datasets, and baselines are provided in the main paper and supplementary material. Upon acceptance, we will release the full code, pretrained models, and instructions for reproducing all reported results, including ablation studies and comparisons. Random seeds and hardware details are also documented to facilitate faithful replication.
@@ A.1 Datasets
3 Degradation Datasets. For both the All-in-One and single-task settings, we follow the evaluation protocols established in prior works Li et al. (2022); Potlapalli et al. (2024); Zamfir et al. (2025), utilizing the following datasets: For image denoising in the single-task setting, we combine the BSD400 Arbelaez et al. (2010) and WED Ma et al. (2016) datasets, and corrupt the images with Gaussian noise at levels $\sigma\in\{15,25,50\}$. BSD400 contains 400 training images, while WED includes 4,744 images. We evaluate the denoising performance on BSD68 Martin et al. (2001) and Urban100 Huang et al. (2015). For single-task deraining, we use Rain100L Yang et al. (2020), which provides 200 clean/rainy image pairs for training and 100 pairs for testing. For single-task dehazing, we adopt the SOTS dataset Li et al. (2018), consisting of 72,135 training images and 500 testing images. Under the All-in-One setting, we train a unified model on the combined set of the aforementioned training datasets for 120 epochs and directly test it across all three restoration tasks.
@@
5 Degradation Datasets. The 5-degradation setting is built upon the 3-degradation setting, with two additional tasks included: deblurring and low-light enhancement. For deblurring, we adopt the GoPro dataset Nah et al. (2017), which contains 2,103 training images and 1,111 testing images. For low-light enhancement, we use the LOL-v1 dataset Wei et al. (2018), consisting of 485 training images and 15 testing images. Note that for the denoising task under the 5-degradation setting, we report results using Gaussian noise with $\sigma=25$. The training takes 130 epochs.
@@
Composited Degradation Datasets. Regarding the composite degradation setting, we use the CDD11 dataset Guo et al. (2024b). CDD11 consists of 1,183 training images for: (i) 4 kinds of single-degradation types: haze (H), low-light (L), rain (R), and snow (S); (ii) 5 kinds of double-degradation types: low-light + haze (l+h), low-light+rain (L+R), low-light + snow (L+S), haze + rain (H+R), and haze + snow (H+S). (iii) 2 kinds of Triple-degradation type: low-light + haze + rain (L+H+R), and low-light + haze + snow (L+H+S). We train our method for 170 epochs (fewer than 200 epochs than MoCE-IR Zamfir et al. (2025)), and we keep all other settings unchanged.
@@
Adverse Weather Removal Datasets. For the deweathering tasks, we follow the experimental setups used in TransWeather Valanarasu et al. (2022) and WGWSNet Zhu et al. (2023), evaluating the performance of our approach on multiple synthetic datasets. We assess the capability of MIRAGE across three challenging tasks: snow removal, rain streak and fog removal, and raindrop removal. The training set, referred to as “AllWeather”, is composed of images from the Snow100K Liu et al. (2018), Raindrop Qian et al. (2018), and Outdoor-Rain Li et al. (2019b) datasets. For testing, we evaluate our model on the following subsets: Snow100K-S (16,611 images), Snow100K-L (16,801 images), Outdoor-Rain (750 images), and Raindrop (249 images). Same as Histoformer Sun et al. (2024), we train MIRAGE on “AllWeather” with 300,000 iterations.
@@
Zero-Shot Underwater Image Enhancement Dataset. For the zero-shot underwater image enhancement setting, we follow the evaluation protocol of DCPT JiaKui et al. (2025) by directly applying our model, trained under the 5-degradation setting, on the UIEB dataset Li et al. (2019a) without any finetuning. UIEB consists of two subsets: 890 raw underwater images with corresponding high-quality reference images, and 60 challenging underwater images. We evaluate our zero-shot performance on the 890-image subset with available reference images.
@@ A.2 Implementation Details
Implementation Details. Our MIRAGE framework is designed to be end-to-end trainable, removing the need for multi-stage optimization of individual components. The architecture adopts a robust 4-level encoder-decoder structure, with a varying number of Mixed Degradation Attention Blocks (MDAB) at each level—specifically $[3,5,5,7]$ from highest to lowest resolution in the Tiny variant. Following prior works Potlapalli et al. (2024); Zamfir et al. (2025), we train the model for 120 epochs with a batch size of 32 in both the 3-Degradation All-in-One and single-task settings. The optimization uses a combination of $L_{1}$ and Fourier loss, optimized with Adam Kingma and Ba (2015) (initial learning rate of $2\text{\times}{10}^{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.999$) and a cosine decay schedule. During training, we apply random cropping to $128{\times}128$ patches, along with horizontal and vertical flipping as data augmentation. All experiments are conducted on a single NVIDIA H200 GPU (140 GB). Memory usage is approximately 42 GB for the Tiny (i.e., MIRAGE -T) model and 56 GB for the Small model (i.e., MIRAGE -S).
@@ A.3 Optimization Objectives
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{1}}+\lambda_{fre}\times\mathcal{L}_{\text{Fourier}}+\lambda_{ctrs}\times\mathcal{L}_{\text{SPD}}.$ (A)
@@
Specifically, we adopt an $\ell_{1}$ loss that adopted in IR tasks Potlapalli et al. (2024); Zamfir et al. (2025); Li et al. (2022); Ren et al. (2026); Cui et al. (2025); Ren et al. (2024); Li et al. (2025), defined as $\mathcal{L}_{\text{1}}=|\hat{x}-x|_{1}$, to enforce pixel-wise similarity between the restored image $\hat{x}$ and the ground-truth image $x$. $\mathcal{L}_{\text{Fourier}}$, as utilized in MoCE-IR Zamfir et al. (2025); Cui et al. (2025), to enhance frequency-domain consistency, the real-valued Fourier loss, is defined as:
$\mathcal{L}_{\text{Fourier}}=\left\|\mathcal{F}_{\text{real}}(\hat{x})-\mathcal{F}_{\text{real}}(x)\right\|_{1}+\left\|\mathcal{F}_{\text{imag}}(\hat{x})-\mathcal{F}_{\text{imag}}(x)\right\|_{1},$ (B)
where $\hat{x}$ and $x$ denote the restored and ground-truth images, respectively. $\mathcal{F}_{\text{real}}(\cdot)$ and $\mathcal{F}_{\text{imag}}(\cdot)$ represent the real and imaginary parts of the 2D real-input FFT (i.e., $\mathrm{rfft2}$). The final loss is computed as the $\ell_{1}$ distance between the real and imaginary components of the predicted and target frequency spectra. Same as MoCE-IR Zamfir et al. (2025), $\lambda_{fre}$ is set to 0.1 throughout our experiments. Meanwhile, the $\mathcal{L}_{\text{SPD}}$ is defined as in Eq. 3-5 of our main manuscript. More ablation studies regarding the proposed $\mathcal{L}_{\text{SPD}}$ are provided in Sec. C.3. The temperature parameter $\tau$ of the proposed $\mathcal{L}_{\text{SPD}}$ is set to 0.1 throughout all the experiments.
Appendix B Preliminaries on SPD-based Feature Statistics
This section provides a brief background on the concepts involved in our cross-layer alignment strategy. The intention is to supply intuitive context—rather than additional derivations—for second-order feature statistics, the SPD structure, and depth-asymmetric representations used in Sec. 4.2.
Second-order feature statistics. Raw activations capture local appearance, but the way channels vary together often reveals more stable information about degradations. For a feature matrix $X\in\mathbb{R}^{C\times N}$, the covariance
$\mathbf{C}=\frac{1}{N-1}(X-\mu)(X-\mu)^{\top}$
summarizes inter-channel relationships. Diagonal entries reflect each channel’s variability, while off-diagonal entries describe redundancy and dependence patterns. These structures differ consistently across layers and degradations (Fig. 4; Appendix Fig. B), making covariance a compact and informative descriptor.
SPD property of covariance matrices. Covariance matrices are symmetric and positive definite by construction and therefore lie in the SPD set. This structure encodes meaningful geometric information: eigenvalues represent correlation strengths, and the matrix as a whole can be interpreted as a “shape” in channel space. Preserving this structure is important—direct Euclidean operations may flatten or distort correlation patterns, an effect also reflected in the collapse observed with Euclidean contrastive learning (Fig. 6).
Representing SPD matrices for comparison. To compare covariance matrices within a contrastive objective, we vectorize $\mathbf{C}$ and apply a learnable projection. This retains second-order relationships while mapping them to an embedding space suitable for contrastive learning. Compared to raw feature vectors, covariance embeddings emphasize structural organization and therefore provide a more stable alignment signal.
Depth-asymmetric representations. Shallow and latent features naturally exhibit different statistical behavior: shallow layers respond strongly to local degradations and show pronounced redundancy, while deeper layers become more decorrelated and semantically aggregated. Their covariance matrices reflect these differences in a consistent way across degradations, making shallow–latent pairs complementary views of the same signal and a natural target for alignment.
Intuition behind SPD-based alignment. Aligning covariance-based SPD embeddings focuses on how channels interact, rather than on individual activation values. This yields supervision that is less sensitive to local noise and more reflective of the underlying representation structure. Encouraging shallow and latent features to share similar second-order statistics stabilizes the shared feature space required for diverse degradations.
Overall, covariance provides a compact view of channel interactions, the SPD structure preserves meaningful second-order relations, and depth-asymmetric covariance patterns naturally motivate the alignment strategy formalized in Sec. 4.2.
Appendix C More Method Details & Supplementary Experiments
C.1 1 Deg. Comparison
@@ Appendix C More Method Details & Supplementary Experiments
DehazeNet - 22.46 .851
MSCNN - 22.06 .908
AODNet - 20.29 .877
EPDN - 22.57 .863
FDGAN - 23.15 .921
AirNet 9M 23.18 .900
PromptIR 36M 31.31 .973
@@
DIDMDN - 23.79 .773
UMR - 32.39 .921
SIRR - 32.37 .926
MSPFN - 33.50 .948
LPNet - 23.15 .921
AirNet 9M 34.90 .977
PromptIR 36M 37.04 .979
@@
DnCNN - 33.89 .930 31.23 .883 27.92 .789
IRCNN - 33.87 .929 31.18 .882 27.88 .790
FFDNet - 33.87 .929 31.21 .882 27.96 .789
BRDNet - 34.10 .929 31.43 .885 28.16 .794
AirNet 9M 34.14 .936 31.48 .893 28.23 .806
PromptIR 36M 34.34 .938 31.71 .897 28.49 .813
PromptIR (Reproduce) 36M 34.15 .934 31.50 .894 28.33 .807
@@
Single-Degradation. In Tab. B, we compare our method against state-of-the-art approaches on single degradation tasks. For dehazing on SOTS dataset, we compare with DehazeNet Cai et al. (2016), MSCNN Ren et al. (2016), AODNet Li et al. (2017), EPDN Qu et al. (2019), FDGAN Dong et al. (2020), and all-in-one methods AirNet Li et al. (2022) and PromptIR Potlapalli et al. (2024). Our 6M parameter model achieves competitive performance (31.46 dB PSNR, 0.977 SSIM), while our 10M model establishes new state-of-the-art results (31.53 dB PSNR, 0.980 SSIM), outperforming the much larger PromptIR (36M parameters). For deraining on Rain100L, we evaluate against DIDMDN Zhang and Patel (2018), UMR Yasarla and Patel (2019), SIRR Wei et al. (2019), MSPFN Jiang et al. (2020), LPNet Gao et al. (2019), AirNet Li et al. (2022), and PromptIR Potlapalli et al. (2024). Our method significantly outperforms all baselines, with our 10M model achieving 38.01 dB PSNR and 0.982 SSIM. For denoising on BSD68, we compare with classical methods DnCNN Zhang et al. (2017a), IRCNN Zhang et al. (2017b), FFDNet Zhang et al. (2018), BRDNet Tian et al. (2020), and recent all-in-one approaches AirNet Li et al. (2022) and PromptIR Potlapalli et al. (2024). Our method consistently outperforms all competitors across different noise levels ($\sigma$=15, 25, 50), demonstrating superior performance with significantly fewer parameters than existing all-in-one methods.
@@
[…]
@@
[…]
@@
[…]
@@
C.2 Details of the Design for the proposed Mixed Backbone.
@@
# patch_embeddingpatch\_embedding: shallow convolutional patch embedding
# refinement_convrefinement\_conv: the refinement block and the final convolution
# spdspd: compute SPD feature
for x{x} in loader: # load a minibatch x with n samples
$F_{\text{shallow}}$ = patch_embeddingpatch\_embedding(x{x}) # Convolutional Patch Embedding
@@
$\mathbf{C}_{s}$, $\mathbf{C}_{l}$= spdspd($F_{\text{shallow}}$), spdspd($F_{\text{latent}}$) # Compurte SPD (Symmetric Positive Definite) manifold features
$\mathbf{z}_{s}$, $\mathbf{z}_{l}$ = proj_normproj\_norm($\mathbf{C}_{s}$), proj_normproj\_norm($\mathbf{C}_{l}$) # Projection and normalize
@@
$\hat{x}$ = refinement_convrefinement\_conv($F_{\text{recon}}$)
[…]
@@
update($f_{en}$, $f_{de}$, patch_embeddingpatch\_embedding, refinement_convrefinement\_conv) # SGD update
@@
def $\mathcal{L}_{\text{SPD}}$(a, b): # SPD Loss
Please refer to Eq.5 of our main manuscript.
@@
C.3 Details of the Proposed SPD Contrastive Learning.
@@
C.4 Ablation Regarding the Optimization Objectives
Tab. C shows that replacing SPD-based contrastive learning with a standard Euclidean-space contrastive loss (w/o SPD) results in a clear performance drop,